### v1

@@ 7 Examples of graph traversals
For all examples considered in this sections, the starting vertex is vertex $1$. We use $d\!=\!2$ for diagonal elements of graph matrices.
Let us compare the traversals implemented by BFS and CCS for the graph $G$ presented at Fig. 2. For this graph, the transformation F of the form (6) is given by the following equations:
@@
Iterations of the traversal, which the algorithm gives for the graph, are presented at Fig. 3. It takes 4 iterations to complete the traversal of the graph using BFS, i.e., Jacobi iterations.
For an CCS iteration, the transformation F of the state vector $x^{(k)}$ is given by the equations (18):
$\left\{\begin{array}[]{l}x_{1}^{(k+1)}=\left(-1+x_{2}^{(k)}\right)d,\\ x_{2}^{(k+1)}=\left(x_{1}^{(k+1)}+x_{3}^{(k)}+x_{6}^{(k)}\right)(-d),\\ x_{3}^{(k+1)}=\left(x_{2}^{(k+1)}+x_{4}^{(k)}+x_{7}^{(k)}\right)(-d),\\ x_{4}^{(k+1)}=\left(x_{3}^{(k+1)}\right)(-d),\\ x_{5}^{(k+1)}=\left(x_{6}^{(k)}\right)(-d),\\ x_{6}^{(k+1)}=\left(x_{2}^{(k+1)}+x_{5}^{(k+1)}+x_{7}^{(k)}\right)(-d),\\ x_{7}^{(k+1)}=\left(x_{3}^{(k+1)}+x_{6}^{(k+1)}+x_{8}^{(k)}\right)(-d),\\ x_{8}^{(k+1)}=\left(x_{7}^{(k+1)}\right)(-d).\\ \end{array}\right.$ (18)18( 18 )
@@
Implementing iterations (18), we obtain a traversal of the graph just for 2 iterations. As we may see, unlike BFS, when performing the first iteration (18) of CCS, a transition will be made along correct chains outgoing from vertex $2$, which is reached from vertex $1$ at the same iteration (Fig. 4).
@@
For BFS, it takes four iterations, which is the maximum possible number of iterations considering all possible starting vertices for a traversal.
@@
Figure 7: Incorrect ordered chain.
@@
With different vertex numbering (Fig. 7) and with the same starting vertex 1, at every CCS iteration, there are no correct chains outgoing from visited vertices. Therefore, to reach all vertices of the graph, both BFS and CSS need to carry out 4 iterations, and the traversal implemented by CCS completely repeat the traversal implemented by BFS (Fig. 8).
@@ 8 Unsigned CCS
The following transformation may be used as the transformation F of the state vector:
$x_{i}^{(k+1)}=\biggl{(}b_{i}+\sum\limits_{l=1}^{i-1}a_{ij}x_{j}^{(k+1)}+\sum\limits_{j=i+1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}d.$ (21)21( 21 )
Let us call as unsigned CCS the algorithm Finding the connected component that uses (21) for F.
Unsigned CCS, implementing the same traversal as CCS, performs computations using only non-negative integer values. The absence of division and possible absence of even multiplication (for $d\!=\!1$) makes it possible to sagnificantly reduce the time required to implement graph traversal.
@@
If we don’t take into account the multiplication by $d$, the unsigned CCS iteration is similar to the BFS iteration, in which each component of the state vector is computed using components with smaller indices already computed at this iteration. But at the same time, the multiplication (division) by $d$ and getting inverse element for addition may be fundamental to apply the graph traversals to solve some discrete analysis problems. For example, implementation of CCS for perturbed graph matrices [13] gives an approach to solve the graph isomorphism problem; it also may be leveraged to perform connectivity test using perturbations of a graph adjacency matrix elements [14] and to computationally solve other problems on graphs.
@@ 9 State vector regularization and vertex masking
Using multiplication in transformations (7) allows us to reduce the time it takes to implement graph traversal. For stability of computations, to prevent overflow or vanishing to zero of the state vector components, after a given number of iterations $M$, the state vector can be regularized:
$x_{i}^{(k)}\to\frac{1}{d^{M}}x_{i}^{(k)}.$ (22)22( 22 )
@@
In addition, without changing the order of the computational complexity of the algorithm, a significant acceleration of implementation of graph traversal can be achieved using vertex masking. Masking a vertex means excluding it from further computations. Masked vertices are the vertices from which new previously unvisited vertices cannot be reached in the course of next iterations of the traversal algorithm. We may both mask the vertices that belong to already founded connected components of a graph and mask already visited vertices of the currently constructed connected component. In the second case, when performing computations at the $(k\!+\!1)$-the iteration (7), vertices that belonged to frontier at the previous $k\!-\!1$ iterations are masked.
@@
For economy of computer memory and to make computions faster, the computations must be carried out using a portrait of adjacency matrix. The portrait contains only non-zero elements of the matrix, i.e., it is the list of its edges. One iteration of the form (6) or (7) has computational complexity of $O(m)$ since we can perform one iteration at one pass through the list of a graph edges.
10 Traversal of a graph
and numbering (labeling) of its vertices
@@ 10 Traversal of a graph
The presence of correct chains, determined by the numbering of the graph vertices, and the choice of the starting vertex are both determine the computational complexity of solving an individual instance of the problem by CCS.
@@
The numbering of vertices is a mandatory parameter of an individual task of implementing graph traversal. Consider the graph for which it is assumed that information is regularly updated with addition of vertices and edges to the graph or with its deletion from it. Optimization of a numbering of graph vertices can be done by numbering them based on their distances from chosen root vertex (starting vertex). With optimal numbering of vertices, to construct a traversal, CCS will have the minimum possible computational complexity for traversal algorithms that is equal to $O(m)$ since we have to do the only one iteration to traverse a graph. To achive this efficiency, we can renumber graph vertices using rechability tree that we obtain after some of the executions of a traversal algorithm for the graph. If the enumeration of the graph vertices is close to the optimal in sense of having correct chains, the computational complexity of CCS also will be of $O(m)$ for the graph.
@@
Note that it may be natural to assign graph vertices numbers, which are reflects distances between objects they are represent. This happens, for example, when numbering objects for transport networks. Randomness in input data is inevitable but it can be minimized. A graph with solely random numbering of vertices, in this case, can be interpreted as a graph of a transport network stored in a database table with randomly generated numbers of records, which may happen in practice only hypothetically with long-term input of data as the network develops. And even for solely random enumerations, there are a lot of instances with correct chains.
@@ 11 The number of iterations required by BFS and CCS
Let us compare the number of iterations required by BFS and CCS to traverse a graph. Since we consider the number of connected components to be of order $O(1)$ with respect to the problem dimension $n$, this value does not change the order of computational complexity of the algorithms. If we do not consider such details of technical realizations as parallelization of computations and masking of the vertices, the computational complexity of a graph traversal algorithm for connected graph is determined by the number of iterations required to visit all its vertices.
Let $N_{\mbox{\tiny BFS}}$ be the number of iterations, which BFS required to complete traversal of a connected graph, and let $N_{\mbox{\tiny CCS}}$ be the number of iterations, which CCS required for the same graph and the same starting vertex. The values $N_{\mbox{\tiny BFS}}$ and $N_{\mbox{\tiny CCS}}$ are determined by the choice of the starting vertex, but $N_{\mbox{\tiny CCS}}$ is also determined by the numbering of vertices in the given graph.
Since both BFS and CCS perform the transitions through edges incident the frontiers’ vertices, we can assert that for any instance of a problem, the number of iterations required to reach all vertices of the connected component for CCS does not exceed the number of such iterations for BFS. At the same time, if, in the course of CCS iterations, we visit a vertex, from which at least one correct chain outgoes, then reachability of vertices from the starting one will be determined in fewer iterations than the algebraic BFS would require for it.
So, since the algorithms stop when we have visited all graph vertices, the following statement is true.
@@ Conclusions
We consider the methods of simple iterations as realizations of graph traversals when solving systems of linear algebraic equations associated with graphs. These systems are systems with modified adjacency matrices of graphs and with the specified right-hand side. There are two different algorithms for graph traversal can be produced by this approach. One of them implemented by iterations of the Jacobi method and the other by iterations of the Gauss-Seidel method. The traversal associated with the Gauss-Seidel method is not equivalent to either the depth-first search or the breadth-first search. For any instance of the problem of finding connected components of a graph, the number of iterations required for such graph traversal does not exceed the number of iterations required for the breadth-first search for the same given starting vertex. For a large number of instances of the problem, fewer iterations will be required for the traversal associated with Gauss-Seidel iterations.

### v2

@@ 7 Examples of graph traversals
For all examples considered in this sections, the starting vertex is vertex $1$. $d\!=\!2$.
Let us compare the traversals implemented by BFS and CCS for the graph $G$ presented at Fig. 2. For this graph, the transformation F of the state vectors $x^{(k)}$ is given by the following equations of the form (6):
@@
Iterations of the traversal, which the iterations gives for the graph, are presented at Fig. 3. It takes 4 iterations to complete the traversal of the graph using BFS, i.e., Jacobi iterations (6).
For an CCS iteration, the transformation F of the state vectors $x^{(k)}$ is given by the equations (18):
$\left\{\begin{array}[]{l}x_{1}^{(k+1)}=\left(-1+x_{2}^{(k)}\right)(-d),\\ x_{2}^{(k+1)}=\left(x_{1}^{(k+1)}+x_{3}^{(k)}+x_{6}^{(k)}\right)(-d),\\ x_{3}^{(k+1)}=\left(x_{2}^{(k+1)}+x_{4}^{(k)}+x_{7}^{(k)}\right)(-d),\\ x_{4}^{(k+1)}=\left(x_{3}^{(k+1)}\right)(-d),\\ x_{5}^{(k+1)}=\left(x_{6}^{(k)}\right)(-d),\\ x_{6}^{(k+1)}=\left(x_{2}^{(k+1)}+x_{5}^{(k+1)}+x_{7}^{(k)}\right)(-d),\\ x_{7}^{(k+1)}=\left(x_{3}^{(k+1)}+x_{6}^{(k+1)}+x_{8}^{(k)}\right)(-d),\\ x_{8}^{(k+1)}=\left(x_{7}^{(k+1)}\right)(-d).\\ \end{array}\right.$ (18)18( 18 )
@@
Implementing iterations (18), we obtain a traversal of the graph just for 2 iterations. As we may see, unlike BFS, when performing the first iteration (18) of CCS, a transition will be made along correct chains outgoing from vertex $2$ which is reached from vertex $1$ at the same iteration (Fig. 4).
@@
For BFS, it takes four iterations which is the maximum possible number of iterations considering all possible starting vertices for a traversal.
@@
Figure 7: Incorrect chain.
@@
With different vertex numbering (Fig. 7), at every CCS iteration, there are no correct chains outgoing from visited vertices. Therefore, to reach all vertices of the graph, both BFS and CSS need to carry out 4 iterations. For this case, the traversal implemented by CCS completely repeat the traversal implemented by BFS (Fig. 8).
@@ 8 Unsigned CCS
The following transformation may be used as the transformation F of the state vector in the algorithm Traversal of a connected component:
$x_{i}^{(k+1)}=\biggl{(}b_{i}+\sum\limits_{j=1}^{i-1}a_{ij}x_{j}^{(k+1)}+\sum\limits_{j=i+1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}d.$ (21)21( 21 )
Let us call as unsigned CCS the algorithm Traversal of a connected com- ponent that uses (21) for F.
Unsigned CCS, implementing the same graph traversal as CCS, performs computations using only non-negative integer values if $d\!\in\!\mathbb{Z}$. The absence of division and possible absence of even multiplication for $d\!=\!1$ allows to sagnificantly reduce the time required to implement a graph traversal.
@@
If we do not take into account the multiplication by $d$, the unsigned CCS iteration is similar to the BFS iteration, in which each component of the state vector is computed using components with smaller indices already computed at this iteration. But at the same time, the multiplication (division) by $d$ and getting inverse element for addition may be fundamental to apply the graph traversals to solve some discrete analysis problems. For example, implementation of CCS for perturbed graph matrices gives an approach to solve the graph isomorphism problem [13]; it also may be leveraged to perform connectivity test using perturbations of a graph adjacency matrix elements [14] and to computationally solve other problems on graphs.
@@ 9 State vector regularization and vertex masking
Using multiplication instead of division in transformation (7) allows us to reduce the time it takes to implement graph traversal. For stability of computations, to prevent overflow or vanishing to zero of the state vector components as machine numbers if $d\!\in\!(0,1)$, after a given number of iterations $M$, the state vector can be regularized:
$x_{i}^{(k)}\leftarrow\frac{1}{d^{M}}\,x_{i}^{(k)}.$ (22)22( 22 )
@@
In addition, without changing the order of computational complexity of CCS, a significant acceleration of graph traversal implementation can be achieved using vertex masking. Masking a vertex means excluding it from further computations. Masked vertices are the vertices from which previously unvisited vertices cannot be reached in the course of next iterations of the traversal algorithm. We can both mask the vertices that belong to already founded connected components of a graph and mask already visited vertices of the currently traversed connected component at an iteration of the Finding all connected components algorithm. In the second case, when performing computations at the $(k+1)$-th iteration (7), vertices that belonged to frontiers obtained at the previous $k\!-\!1$ iterations of the algorithm Traversal of a connected component are masked.
@@
For economy of computer memory and to make computions faster, the computations must be carried out using a portrait of adjacency matrix. The portrait contains only non-zero elements of the matrix, i.e., it is the list of graph’s edges. Using the portrait, implementation of one iteration of the form (6) or (7) has computational complexity of $O(m)$ since we can perform one iteration at one pass through the list of a graph edges.
10 Traversal of a graph and numbering of its vertices
@@
The presence of correct chains in a graph is determined by the numbering of the graph vertices. So this numbering and the choice of the starting vertex are both determine the computational complexity of a graph traversal by CCS. The numbering of a graph’s vertices is a mandatory parameter of an instance of the problem on graphs.
@@
Consider the graph for which it is assumed that information is regularly updated with addition of vertices and edges to the graph or with its deletion from it. Optimization of a numbering of graph vertices can be done by numbering them based on their distances from chosen vertex that is starting vertex. With optimal numbering of vertices, for which there are correct chains that connects starting vertex with any other vertex, to perform a traversal, CCS will have the minimum possible computational complexity for traversal algorithms that is equal to $O(m)$ since we have to do the only one iteration to traverse a graph. To achive this efficiency, we can renumber graph vertices using rechability tree that we obtain after previous executions of a traversal algorithm for the graph of a problem which is modified. If the enumeration of the graph vertices is close to the optimal in sense of having correct chains, the computational complexity of CCS also will be of $O(m)$ for the graph.
@@
Note that it may be natural to assign graph vertices numbers which are reflects distances between objects they are represent. This happens, for example, when numbering objects for transport networks. Randomness in such input data as numbering of graph’s vertices is inevitable in applications but it can be minimized. A graph with solely random numbering of vertices can be interpreted as a graph of a transport network stored in a database table with randomly generated numbers of records, which may happen in practice only hypothetically with long-term input of data as the network develops. And even for solely random enumerations, there are a lot of instances with correct chains.
@@ 11 The number of iterations required by BFS and CCS
to traverse a connected graph
Let us compare the number of iterations required by BFS and CCS to traverse a graph. If we do not consider such details of realizations of the traversals as parallelization of computations and masking of the vertices, the computational complexity of a graph traversal algorithm for connected graph is determined by the number of iterations required to visit all its vertices.
Let $N_{\mbox{\tiny BFS}}$ be the number of iterations which BFS required to complete traversal of a connected graph, and let $N_{\mbox{\tiny CCS}}$ be the number of iterations which CCS required for the same graph and the same starting vertex. The values $N_{\mbox{\tiny BFS}}$ and $N_{\mbox{\tiny CCS}}$ are determined by the choice of the starting vertex, but $N_{\mbox{\tiny CCS}}$ is also determined by the numbering of vertices in the given graph.
Since both BFS and CCS perform the transitions through edges incident the frontiers’ vertices, we can assert that for any graph $N_{\mbox{\tiny CCS}}\!\leq\!N_{\mbox{\tiny BFS}}$. At the same time, if, in the course of CCS iterations, we visit at least one vertex, from which at least one correct chain outgoes, then reachability of vertices from the starting one will be determined in fewer iterations than the algebraic BFS would require for it, i.e., $N_{\mbox{\tiny CCS}}\!<\!N_{\mbox{\tiny BFS}}$ in this case. So the following statement is true.
@@ Conclusions
We consider the methods of simple iterations for solving systems of linear algebraic equations associated with graphs as realizations of graph traversals. These systems are systems with modified adjacency matrices of graphs and with the specified right-hand side. There are two different algorithms for graph traversal can be produced by this approach. One of them implemented by iterations of the Jacobi method and the other by iterations of the Gauss-Seidel method. The traversal associated with the Jacobi method is equivalent to the breadth first search, and the traversal associated with the Gauss-Seidel method is not equivalent to either the depth-first search or the breadth-first search. For any instance of the problem of finding connected components of a graph, the number of iterations required for the presented graph traversal does not exceed the number of iterations required for the breadth-first search for the same given starting vertex. For a large number of instances of the problem, fewer iterations will be required for the traversal associated with Gauss-Seidel iterations.